# Performance bounds for priority-based stochastic coflow scheduling

### v1

@@ Abstract
We consider the coflow scheduling problem in the non-clairvoyant setting, assuming that flow sizes are realized on-line according to given probability distributions. The goal is to minimize the weighted average completion time of coflows in expectation. We first obtain inequalities for this problem that are valid for all non-anticipative order-based rate-allocation policies and define a polyhedral relaxation of the performance space of such scheduling policies. This relaxation is used to analyze the performance of a simple priority policy in which the priority order is computed by $\mathtt{Sincronia}$from expected flow sizes instead of their unknown actual values. We establish a bound on the approximation ratio of this priority policy with respect to the optimal priority policy for arbitrary probability distributions of flow sizes (with finite first and second moments). Tighter upper bounds are obtained for some specific distributions. Extensive numerical results suggest that performance of the proposed policy is much better than the upper bound.
@@ 1 Introduction
Modern data centres are increasingly used to run data-parallel computing applications such as MapReduce, Hadoop and Spark [13, 33]. These applications alternate between computation stages and communication stages, during which the application’s tasks exchange intermediate results using the data centre network. Each application may generate hundreds or thousands of concurrent data transfers which have to share the resources of the datacenter network not only between themselves, but also with the data flows of the other applications. Chowdhury and Stoica showed in [10] that by allocating rates to all these data transfers in some appropriate way, it was possible to significantly reduce the average time spent in communication by data-parallel applications. They introduced the coflow scheduling problem, a coflow being defined as a collection of flows generated by the same parallel applicaton and its completion time being determined by the completion time of the last flow in the collection (that is, it corresponds to the time at which the next computation stage can start). The coflow scheduling problem is then to allocate rates to flows over time so that the average coflow completion time (CCT) is minimized.
@@
Even in the clairvoyant offline setting, in which all coflows to be scheduled are initially present in the system and the precise volume of each and every flow is known, the problem is known to be NP-hard [12]. This follows from the fact that it can be seen as a generalization of the concurrent open-shop scheduling problem in which resources are coupled, as transmitting a flow requires capacity on both the input and output ports. Furthermore, the coflow scheduling problem was shown to be inapproximable below a factor of $2$ [5, 25] in this setting. Several heuristics and approximation algorithms for clairvoyant coflow scheduling have been proposed in the last decade, see e.g. [12, 8, 31, 32, 9, 19, 28, 35]. At the present time, the state-of-the-art algorithm for coflow scheduling in the clairvoyant setting is $\mathtt{Sincronia}$ [3], which was proposed by Agarwal et al. in 2018 and provides an approximation ratio of $4$, the best known one (see also [27] and [4] for closely-related references that were published roughly concurrently with [3]) . In addition, an attractive feature of $\mathtt{Sincronia}$is that it relies on priority mechanisms which are supported by most transport layers. It is worth mentioning that $\mathtt{Sincronia}$is not restricted to the offline setting but can also efficiently handle online coflow arrivals.
@@
Although $\mathtt{Sincronia}$provides a fairly efficient solution to the coflow scheduling problem, it requires a priori knowledge of the flow volumes, which is difficult to obtain in practice. Therefore, some authors have investigated the non-clairvoyant setting, where on a coflow arrival, only the number of flows and their input/output ports are revealed to the coflow scheduler, while their volumes remains unknown [11, 36, 15, 18, 34, 14, 6]. However, none of the proposed algorithms comes with performance guarantees, with the noticeable exception of the $\mathtt{BlindFlow}$algorithm which was shown in [6] to be $8\,p$-approximate, where $p$ is the maximum number of flows that any coflow can have. Given the orders of magnitude of $p$ in practice, this performance guarantee is of course much weaker than that provided by $\mathtt{Sincronia}$in the clairvoyant setting.
@@
It could be argued however that assuming absolutely no prior information on flow volumes, as was done in the above works on the non-clairvoyant setting, is excessively pessimistic. Indeed, historical data on previous executions of data-parallel applications could be exploited to get some insights on what the flow volumes could be. For instance, machine learning (ML) algorithms could be used to obtain predictions on flow volumes from historical data. The improvement of online scheduling algorithms via unreliable ML predictions is a recent line of research which has attracted a lot of studies in the last few years (see e.g.  [23, 16]) and which was recently applied to non-clairvoyant coflow scheduling [7]. Another relevant possibility, which is explored in the present paper, is to assume that the unknown flow volumes are on-line realizations of known probability distributions, which can be estimated from flow sizes data collected during previous executions.
@@
Stochastic coflow scheduling has received little attention up to now. To the extent of our knowledge, the only work that has addressed the coflow scheduling problem in the stochastic setting is [19], in which the authors study the non-preemptive coflow scheduling problem. The objective is to minimize the weighted expected completion time of coflows. The proposed approach uses a time-indexed linear programming relaxation first introduced in [29], and uses its solution to come up with a feasible schedule. The proposed algorithm is shown to achieve an approximation ratio of $(1+2\log L)(1+\sqrt{L}CV_{max})(1+LCV_{max})(1+CV_{max}^{2})$, where $L$ is the number of ports and $CV_{max}$ is an upper bound on the coefficient of variation of processing times [^1] The processing time of a coflow on a port is the total time required for transmitting all of its constituent flows on that port in isolation.. We, on the other hand, consider the preemptive case.
@@ 1.1 Contributions
We consider the non-clairvoyant setting in which the size of each flow is unknown to the coflow scheduler and is revealed to it only upon flow completion. Although flow sizes are not known in advance, we assume that they are drawn from known probability distributions. It is shown that for any order-based rate-allocation policy, the vector of expected completion times satisfies certain inequalities, from which a polyhedral relaxation of the performance space of such coflow scheduling policies is obtained. We use this polyhedral relaxation to analyze the performance of a priority policy in which the priority order is computed with $\mathtt{Sincronia}$from expected flow sizes. Specifically, its is shown that the approximation ratio of this policy with respect to the optimal priority policy is upper bounded by $4(1+\sqrt{L}\sigma_{max}/\mu_{min})$ for any distribution of flow sizes with finite first and second moments, where $\mu_{min}$ is a lower bound on the expected processing times and $\sigma_{max}$ is an upper bound on their standard deviations. Tighter bounds can be obtained if we have information on the distribution of processing times. For instance, we show that for normally distributed processing times, the approximation ratio of the $\mathtt{Sincronia}$-based priority policy is upper bounded by $4(1+\sqrt{2\log L}CV_{max})$, where, as before, $CV_{max}$ is an upper bound on the coefficient of variation of processing times. We also obtain asymptotic bounds for Pareto-distributed processing times.
@@ 2 Stochastic coflow scheduling
We consider a set ${\mathcal{C}}=\{1,2,\ldots,n\}$ of coflows, which are all present in the system at time 00. Each coflow $k$ is a set $F_{k}$ of flows, where flow $j$ of coflow $k$ is characterized by its source node, its destination node and its volume $V^{k,j}$. Each coflow $k$ may also be assigned a weight $w_{k}$ (default weight is $1$). As for the network, we assume the big switch model [^2] This model abstracts out the datacenter network fabric as one big switch interconnecting servers. The underlying assumption is that the fabric core can sustain 100% throughput and only the ingress and egress ports are potential congestion points. [12] and we let $b_{\ell}$ denote the capacity of port $\ell\in{\mathcal{L}}$. We let $F_{k,\ell}$ be the set of flows of coflow $k$ that use port $\ell$ either as ingress or egress port, and denote by $C_{k}$ the completion time of coflow $k$, that is, the time at which its last flow finishes.
@@
A rate allocation policy $r$ specifies the rate allocated to each flow at any given time in such a way that constraint (1) is satisfied. We shall specifically consider work-conserving policies which allocate a constant rate to flows in between decision epochs, where decision epochs correspond to time 00 and to flow completion times. If $t$ is a decision epoch, the decisions of such a policy may only depend on the ”past up to time $t$”, which is given by the sets of data transfers already finished or being performed at $t$, as well as on the a priori knowledge of the input data of the problem. This includes of course the distributional information about remaining sizes of unfinished flows. In other words, admissible policies must be nonanticipative in the sense that scheduling decisions cannot depend on future information, such as the actual volumes of unfinished flows. A simple example of a work-conserving and nonanticipative scheduling policy is the round-robin (RR) policy which allocates rate $r^{k,j}(t)=\min\left\{b_{i}/n_{i}(t),b_{o}/n_{o}(t)\right\}$ to flow $j\in F_{k}$ at time $t$, where $i$ and $o$ are the input and output ports of this flow, respectively, and $n_{\ell}(t)$ is the number of ongoing flows on port $\ell$ at time $t$.
Among work-conserving and nonanticipative rate-allocation policies, we restrict ourselves to order-based policies. Such policies define a partial order on the set of flows such that one flow precedes another one if they have either the same source or the same destination. Under order-based policies, no single port is transmitting multiple flows at the same time. The analysis of order-based policies is much simpler as it is possible to leverage a well-known reduction to the Concurrent Open Shop problem [17]. Moreover, a key advantage of these policies is that they are easily implementable by leveraging priority mechanisms supported by most transport layers. An important property of order-based policies is that the completion order of the flows (and hence of coflows) on all ports is fixed and independent of flow sizes. In the following, we let ${\mathcal{R}}$ be the set of order-based policies. Note that the RR policy does not belong to ${\mathcal{R}}$ as multiple flows may be served in parallel. However, ${\mathcal{R}}$ contains all priority policies which schedule coflows in some priority order $\pi$. Such a policy allocates a positive rate to a flow of coflow $\pi(k^{\prime})$ if and only if no flows of a coflow $\pi(k)$ such that $k<k^{\prime}$ is available for transmission on the same input/output port.
@@
[…]
@@ 3 Sincronia with expected processing times as inputs
In this section, we establish a generic bound on the approximation ratio of $\mathtt{Sincronia}$for problem (P2) when ran with expected processing times as inputs. Following an approach proposed in [21], we first show in Section 3.1 that for any order-based policy $r$, the vector of expected completion times $\mathbb{E}\left[C_{k}(r)\right]$ satisfies certain inequalities. The latter inequalities yield a polyhedral relaxation of the performance space of order-based scheduling policies, as discussed in Section 3.2. It turns out that this polyhedral relaxation is exactly the same as the one used in $\mathtt{Sincronia}$in the clairvoyant setting, except that the actual processing times are replaced by their expected values. In Section 3.3, we briefly describes how $\mathtt{Sincronia}$works and propose a priority policy for stochastic coflow scheduling in which rates are allocated to coflows according to the priority order computed by $\mathtt{Sincronia}$from expected processing times. We obtain performance guarantees for this policy in Section 3.4.
@@ 3.1 Stochastic Parallel Inequalities
The so-called parallel inequalities are inequalities on the completion times of jobs which are valid for any feasible schedule on $m$ machines [24, 26]. These inequalities were extended in [21] to stochastic parallel-machine scheduling. We adapt the proof of Theorem 3.1 in [21] to show that the same inequalities are also valid for the stochastic coflow scheduling problem.
@@
for all $l\in{\mathcal{L}}$ and $A\subseteq{\mathcal{C}}$.
@@ 3.3 Sincronia-based priority policy
In order to schedule a set ${\mathcal{C}}$ of coflows with known processing times $p_{\ell,k}$, $\mathtt{Sincronia}$ first computes a priority order over coflows, that is, a permutation $\pi$ of ${\mathcal{C}}$. If $\pi(k)<\pi(k^{\prime})$, then the flows of coflow $k$ are transmitted in the data centre network with strict priority over the flows of coflow $k^{\prime}$. $\mathtt{Sincronia}$then uses a preemptive greedy rate allocation which assigns to a single flow the full rate of its incoming and outgoing ports at any given point of time and preserves this ordering, that is, it is guaranteed that coflows will be completed in the priority order (that is, $C_{k}\leq C_{k^{\prime}}$ if $\pi(k)<\pi(k^{\prime}))$.
It turns out that problem (LP-Primal) is exactly the same as the one used in $\mathtt{Sincronia}$to compute the priority order $\pi$ [3], except that the actual processing times $p_{\ell,k}$ have been replaced by their expected values $\mu_{\ell,k}$. A natural approach to obtain an admissible scheduling policy in the stochastic setting is therefore to compute a priority order $\pi$ using $\mathtt{Sincronia}$with expected processing times $\mu_{\ell,k}$, instead of the unknown actual processing times $P_{\ell,k}$. The greedy rate allocation of $\mathtt{Sincronia}$then gives an admissible rate-allocation policy for problem (P2).
Before establishing performance guarantees for the above $\mathtt{Sincronia}$-based priority policy, we first briefly describe how $\mathtt{Sincronia}$computes the priority order $\pi$ from expected processing times $\mu_{\ell,k}$. $\mathtt{Sincronia}$uses a primal-dual algorithm in which the primal problem is (LP-Primal), and the dual problem is as follows
@@
The $\mathtt{Sincronia}$primal-dual algorithm is shown in Algorithm 1. The dual variables $y_{\ell,A}$ are initialized to 00 and the set $A$ of unscheduled coflows is initially set to ${\mathcal{C}}$. The algorithm assigns priorities to coflows in the order of increasing priority. At each iteration $t$, the algorithm determines the bottleneck port $b$, that is, the port $\ell$ for which the total expected completion time $\sum_{k\in A}\mu_{\ell,k}$ is maximum. It then determines the coflow $k^{*}$ with the largest weighted expected processing time on the bottleneck. Coflow $k^{*}$ is assigned priority $t$ with $\pi(k^{*})=t$ and it is removed from the set of unscheduled coflows $A$. Before that, the dual variable $y_{b,A}$ and the primal variable $C_{k^{*}}$ are set to their final values and the weights of the coflows using the bottleneck port are updated. As discussed in in [3] , the complexity of Algorithm 1 is $O(n^{2})$.
Algorithm 1 $\mathtt{Sincronia}$primal-dual algorithm with expected processing times $\mu_{\ell,k}$ as inputs
1:$y_{\ell,A}\leftarrow 0$ for all $\ell\in{\mathcal{L}}$ and $A\subseteq{\mathcal{C}}$.
2:$A\leftarrow{\mathcal{C}}$
@@
4:     $b\leftarrow\mbox{argmax}_{\ell\in{\mathcal{L}}}\sum_{k\in A}\mu_{\ell,k}$triangleright Bottleneck port
5:     $k^{*}\leftarrow\mbox{argmin}_{k\in A}\left(w_{k}/\mu_{b,k}\right)$triangleright Largest weighted proc. time
6:     $C_{k^{*}}\leftarrow\sum_{k\in A}\mu_{b,k}$triangleright Set primal variable
7:     $y_{b,A}\leftarrow w_{k^{*}}/\mu_{b,{k^{*}}}$triangleright Set dual variable
8:     $w_{k}\leftarrow w_{k}-w_{k^{*}}\mu_{b,k}/\mu_{b,{k^{*}}}$ for all $k\in A$triangleright Update weights
9:     $\pi(t)\leftarrow k^{*}$triangleright Set priority of coflow $k^{*}$
10:     $A\leftarrow A\setminus\{k^{*}\}$triangleright Remove $k^{*}$ from unscheduled coflows
@@
The approximation ratio of Algorithm 1 is established in [3] and formally stated in Theorem 2 below.

### v2

@@ Abstract
We consider the coflow scheduling problem in the non-clairvoyant setting, assuming that flow sizes are realized on-line according to given probability distributions. The goal is to minimize the weighted average completion time of coflows in expectation. We first obtain inequalities for this problem that are valid for all non-anticipative order-based rate-allocation policies and define a polyhedral relaxation of the performance space of such scheduling policies. This relaxation is used to analyze the performance of a simple priority policy in which the priority order is computed by $\mathtt{Sincronia}\$from expected flow sizes instead of their unknown actual values. We establish a bound on the approximation ratio of this priority policy with respect to the optimal priority policy for arbitrary probability distributions of flow sizes (with finite first and second moments). Tighter upper bounds are obtained for some specific distributions. Extensive numerical results suggest that performance of the proposed policy is much better than the upper bound.
@@ 1 Introduction
Modern data centres are increasingly used to run data-parallel computing applications such as MapReduce, Hadoop and Spark [dean2004mapreduce, zaharia2010spark]. These applications alternate between computation stages and communication stages, during which the application’s tasks exchange intermediate results using the data centre network. Each application may generate hundreds or thousands of concurrent data transfers which have to share the resources of the datacenter network not only between themselves, but also with the data flows of the other applications. Chowdhury and Stoica showed in [ChowdhuryHotNet2012] that by allocating rates to all these data transfers in some appropriate way, it was possible to significantly reduce the average time spent in communication by data-parallel applications. They introduced the coflow scheduling problem, a coflow being defined as a collection of flows generated by the same parallel applicaton and its completion time being determined by the completion time of the last flow in the collection (that is, it corresponds to the time at which the next computation stage can start). The coflow scheduling problem is then to allocate rates to flows over time so that the average coflow completion time (CCT) is minimized.
@@
Even in the clairvoyant offline setting, in which all coflows to be scheduled are initially present in the system and the precise volume of each and every flow is known, the problem is known to be NP-hard [Chowdhury2014]. This follows from the fact that it can be seen as a generalization of the concurrent open-shop scheduling problem in which resources are coupled, as transmitting a flow requires capacity on both the input and output ports. Furthermore, the coflow scheduling problem was shown to be inapproximable below a factor of $2$ [Bansal2010, Sachdeva2013] in this setting. Several heuristics and approximation algorithms for clairvoyant coflow scheduling have been proposed in the last decade, see e.g. [Chowdhury2014, Chen2016, Wang2017, Wang2019, chowdhury2019, Mao2018, Shi2021, Zhang2021]. At the present time, the state-of-the-art algorithm for coflow scheduling in the clairvoyant setting is $\mathtt{Sincronia}\$ [Agarwal2018], which was proposed by Agarwal et al. in 2018 and provides an approximation ratio of $4$, the best known one (see also [Shafiee2018] and [Ahmadi2020] for closely-related references that were published roughly concurrently with [Agarwal2018]) . In addition, an attractive feature of $\mathtt{Sincronia}\$is that it relies on priority mechanisms which are supported by most transport layers. It is worth mentioning that $\mathtt{Sincronia}\$is not restricted to the offline setting but can also efficiently handle online coflow arrivals.
@@
Although $\mathtt{Sincronia}\$provides a fairly efficient solution to the coflow scheduling problem, it requires a priori knowledge of the flow volumes, which is difficult to obtain in practice. Therefore, some authors have investigated the non-clairvoyant setting, where on a coflow arrival, only the number of flows and their input/output ports are revealed to the coflow scheduler, while their volumes remains unknown [Chow2015, Zhang2018, Gao2016, Liu2020, Zhang2016, Dogar2014, Bhimaraju2020]. However, none of the proposed algorithms comes with performance guarantees, with the noticeable exception of the $\mathtt{BlindFlow}\$algorithm which was shown in [Bhimaraju2020] to be 8p8\,p-approximate, where $p$ is the maximum number of flows that any coflow can have. Given the orders of magnitude of $p$ in practice, this performance guarantee is of course much weaker than that provided by $\mathtt{Sincronia}\$in the clairvoyant setting.
@@
It could be argued however that assuming absolutely no prior information on flow volumes, as was done in the above works on the non-clairvoyant setting, is excessively pessimistic. Indeed, historical data on previous executions of data-parallel applications could be exploited to get some insights on what the flow volumes could be. For instance, machine learning (ML) algorithms could be used to obtain predictions on flow volumes from historical data. The improvement of online scheduling algorithms via unreliable ML predictions is a recent line of research which has attracted a lot of studies in the last few years (see e.g.  [Purohit2018, Im2022]) and which was recently applied to non-clairvoyant coflow scheduling [Brun2023]. Another relevant possibility, which is explored in the present paper, is to assume that the unknown flow volumes are on-line realizations of known probability distributions, which can be estimated from flow sizes data collected during previous executions.
@@
Stochastic coflow scheduling has received little attention up to now. To the extent of our knowledge, the only work that has addressed the coflow scheduling problem in the stochastic setting is [Mao2018], in which the authors study the non-preemptive coflow scheduling problem. The objective is to minimize the weighted expected completion time of coflows. The proposed approach uses a time-indexed linear programming relaxation first introduced in [Skutella2016], and uses its solution to come up with a feasible schedule. The proposed algorithm is shown to achieve an approximation ratio of $(1+2\log L)(1+\sqrt{L}\eta_{max})(1+L\eta_{max})(1+\eta_{max}^{2})$, where $L$ is the number of ports and $\eta_{max}$ is an upper bound on the coefficient of variation of processing times [^1] The processing time of a coflow on a port is the total time required for transmitting all of its constituent flows on that port in isolation.. We, on the other hand, consider the preemptive case.
Our work relies on the assumption that the scheduler has knowledge of the statistics of flow volumes of the coflows. The scheduling decisions are then based purely on the mean flow volumes. A complementary line of work to ours is when the statistics are not known to the scheduler but it can obtain them by scheduling a small percentage of flows from every coflow. Once an estimation is obtained, the scheduler can apply a policy of its choice based on this estimation. This line of work was investigated in [JHL2022] in which they proposed $\mathtt{Philae}$, an algorithm to estimate the statistics from executing a small subset of coflows. Six different scheduling policies were compared in $\mathtt{Philae}$, couple of them based on Aalo [CS2015] while some of them were new. The performance of these policies was compared numerically and no guarantees in terms of approximation ratio was provided. As mentioned in [JHL2022], $\mathtt{Sincronia}\$is orthogonal to $\mathtt{Philae}$ $\ldots$”. $\mathtt{Philae}$’s estimation method can be used in conjunction with $\mathtt{Sincronia}\$as the scheduling policy. Our aim is to provide an approximation ratio for $\mathtt{Sincronia}\$when it computes the priority order based on mean flow volumes whereas the work of [JHL2022] proposes a method to obtain estimates of mean flow volumes in real-time but does not focus on the computation of the approximation ratio. A related sampling and estimation technique was proposed in [JHLD2023] but for cluster job scheduling.
@@ 1.1 Contributions
We consider the non-clairvoyant setting in which the size of each flow is unknown to the coflow scheduler and is revealed to it only upon flow completion. Although flow sizes are not known in advance, we assume that they are drawn from known probability distributions. It is shown that for any order-based rate-allocation policy, the vector of expected completion times satisfies certain inequalities, from which a polyhedral relaxation of the performance space of such coflow scheduling policies is obtained. We use this polyhedral relaxation to analyze the performance of a priority policy in which the priority order is computed with $\mathtt{Sincronia}\$from expected flow sizes. Specifically, its is shown that the approximation ratio of this policy with respect to the optimal priority policy is upper bounded by $4(1+\sqrt{L}\sigma_{max}/\mu_{min})$ for any distribution of flow sizes with finite first and second moments, where $\mu_{min}$ is a lower bound on the expected processing times and $\sigma_{max}$ is an upper bound on their standard deviations. Tighter bounds can be obtained if we have information on the distribution of processing times. For instance, we show that for normally distributed processing times, the approximation ratio of the $\mathtt{Sincronia}$-based priority policy is upper bounded by $4(1+\sqrt{2\log L}\eta_{max})$, where, as before, $\eta_{max}$ is an upper bound on the coefficient of variation of processing times. We also obtain asymptotic bounds for Pareto-distributed processing times.
@@ 2 Stochastic coflow scheduling
We consider a set ${\mathcal{C}}=\{1,2,\ldots,n\}$ of coflows, which are all present in the system at time 0. Each coflow $k$ is a set $F_{k}$ of flows, where flow $j$ of coflow $k$ is characterized by its source node, its destination node and its volume $V^{k,j}$. Each coflow $k$ may also be assigned a weight $w_{k}$ (default weight is $1$). As for the network, we assume the big switch model [^2] This model abstracts out the datacenter network fabric as one big switch interconnecting servers. The underlying assumption is that the fabric core can sustain 100% throughput and only the ingress and egress ports are potential congestion points. [Chowdhury2014] and we let $b_{\ell}$ denote the capacity of port $\ell\in{\mathcal{L}}$. We let $F_{k,\ell}$ be the set of flows of coflow $k$ that use port $\ell$ either as ingress or egress port, and denote by $C_{k}$ the completion time of coflow $k$, that is, the time at which its last flow finishes.
@@
A rate allocation policy $r$ specifies the rate allocated to each flow at any given time in such a way that constraint (1) is satisfied. We shall specifically consider work-conserving policies which allocate a constant rate to flows in between decision epochs, where decision epochs correspond to time 0 and to flow completion times. If $t$ is a decision epoch, the decisions of such a policy may only depend on the ”past up to time tt”, which is given by the sets of data transfers already finished or being performed at $t$, as well as on the a priori knowledge of the input data of the problem. This includes of course the distributional information about remaining sizes of unfinished flows. In other words, admissible policies must be nonanticipative in the sense that scheduling decisions cannot depend on future information, such as the actual volumes of unfinished flows. A simple example of a work-conserving and nonanticipative scheduling policy is the round-robin (RR) policy which allocates rate $r^{k,j}(t)=\min\left\{b_{i}/n_{i}(t),b_{o}/n_{o}(t)\right\}$ to flow $j\in F_{k}$ at time $t$, where $i$ and $o$ are the input and output ports of this flow, respectively, and $n_{\ell}(t)$ is the number of ongoing flows on port $\ell$ at time $t$.
Among work-conserving and nonanticipative rate-allocation policies, we restrict ourselves to order-based policies. Such policies define a partial order on the set of flows such that one flow precedes another one if they have either the same source or the same destination. Under order-based policies, no single port is transmitting multiple flows at the same time. The analysis of order-based policies is much simpler as it is possible to leverage a well-known reduction to the Concurrent Open Shop problem [Khuller2018]. Moreover, a key advantage of these policies is that they are easily implementable by leveraging priority mechanisms supported by most transport layers. An important property of order-based policies is that the completion order of the flows (and hence of coflows) on all ports is fixed and independent of flow sizes. In the following, we let $\mathcal{R}$} be the set of order-based policies. Note that the RR policy does not belong to $\mathcal{R}$} as multiple flows may be served in parallel. However, $\mathcal{R}$} contains all priority policies which schedule coflows in some priority order $\pi$. Such a policy allocates a positive rate to a flow of coflow $\pi(k^{\prime})$ if and only if no flows of a coflow $\pi(k)$ such that $k<k^{\prime}$ is available for transmission on the same input/output port.
@@
$\underset{r}{\mathrm{min}}\enskip$ $\sum_{k\in\mathcal{C}}w_{k}\mathbb{E}\left[C_{k}(r)\right]$ (P2)
@@ 3 Sincronia with expected processing times as inputs
In this section, we establish a generic bound on the approximation ratio of $\mathtt{Sincronia}\$for problem (P2) when ran with expected processing times as inputs. Following an approach proposed in [Mohring99], we first show in Section 3.1 that for any order-based policy $r$, the vector of expected completion times $\mathbb{E}\left[C_{k}(r)\right]$ satisfies certain inequalities. The latter inequalities yield a polyhedral relaxation of the performance space of order-based scheduling policies, as discussed in Section 3.2. It turns out that this polyhedral relaxation is exactly the same as the one used in $\mathtt{Sincronia}\$in the clairvoyant setting, except that the actual processing times are replaced by their expected values. In Section 3.3, we briefly describes how $\mathtt{Sincronia}\$works and propose a priority policy for stochastic coflow scheduling in which rates are allocated to coflows according to the priority order computed by $\mathtt{Sincronia}\$from expected processing times. We obtain performance guarantees for this policy in Section 3.4.
@@ 3.1 Stochastic Parallel Inequalities
The so-called parallel inequalities are inequalities on the completion times of jobs which are valid for any feasible schedule on $m$ machines [Queyranne1993, Schulz1996SchedulingTM]. These inequalities were extended in [Mohring99] to stochastic parallel-machine scheduling. We adapt the proof of Theorem 3.1 in [Mohring99] to show that the same inequalities are also valid for the stochastic coflow scheduling problem.
@@
for all $\ell\in{\mathcal{L}}$ and $A\subseteq{\mathcal{C}}$.
@@ 3.3 Sincronia-based priority policy
In order to schedule a set $\mathcal{C}$} of coflows with known processing times $p_{\ell,k}$, $\mathtt{Sincronia}\$first computes a priority order over coflows, that is, a permutation $\pi$ of $\mathcal{C}$}. If $k<k^{\prime}$, then the flows of coflow $\pi(k)$ are transmitted in the data centre network with strict priority over the flows of coflow $\pi(k^{\prime})$. $\mathtt{Sincronia}\$then uses a preemptive greedy rate allocation which assigns to a single flow the full rate of its incoming and outgoing ports at any given point of time and preserves this ordering, that is, it is guaranteed that coflows will be completed in the priority order (that is, $C_{\pi(k)}\leq C_{\pi(k^{\prime})}$ if $k<k^{\prime})$.
It turns out that problem (LP-Primal) is exactly the same as the one used in $\mathtt{Sincronia}\$to compute the priority order $\pi$ [Agarwal2018], except that the actual processing times $p_{\ell,k}$ have been replaced by their expected values $\mu_{\ell,k}$. A natural approach to obtain an admissible scheduling policy in the stochastic setting is therefore to compute a priority order $\pi$ using $\mathtt{Sincronia}\$with expected processing times $\mu_{\ell,k}$, instead of the unknown actual processing times $P_{\ell,k}$. The greedy rate allocation of $\mathtt{Sincronia}\$then gives an admissible rate-allocation policy for problem (P2).
Before establishing performance guarantees for the above $\mathtt{Sincronia}$-based priority policy, we first briefly describe how $\mathtt{Sincronia}\$computes the priority order $\pi$ from expected processing times $\mu_{\ell,k}$. $\mathtt{Sincronia}\$uses a primal-dual algorithm in which the primal problem is (LP-Primal), and the dual problem is as follows
@@
The $\mathtt{Sincronia}\$primal-dual algorithm is shown in Algorithm 1. The dual variables $y_{\ell,A}$ are initialized to 0 and the set $A$ of unscheduled coflows is initially set to $\mathcal{C}$}. The algorithm assigns priorities to coflows in the order of increasing priority. At each iteration $t$, the algorithm determines the bottleneck port $b$, that is, the port $\ell$ for which the total expected completion time $\sum_{k\in A}\mu_{\ell,k}$ is maximum. It then determines the coflow $k^{*}$ with the largest weighted expected processing time on the bottleneck. Coflow $k^{*}$ is assigned priority $t$ with $\pi(t)=k^{*}$ and it is removed from the set of unscheduled coflows $A$. Before that, the dual variable $y_{b,A}$ and the primal variable $C_{k^{*}}$ are set to their final values and the weights of the coflows using the bottleneck port are updated. As discussed in in [Agarwal2018] , the complexity of Algorithm 1 is $O(n^{2})$.
Algorithm 1 $\mathtt{Sincronia}\$primal-dual algorithm with expected processing times $\mu_{\ell,k}$ as inputs
[…]
@@
4:  $b\leftarrow\mbox{argmax}_{\ell\in{\mathcal{L}}}\sum_{k\in A}\mu_{\ell,k}$ $\triangleright$ Bottleneck port
5:  $k^{*}\leftarrow\mbox{argmin}_{k\in A}\left(w_{k}/\mu_{b,k}\right)$ $\triangleright$ Largest weighted proc. time
6:  $C_{k^{*}}\leftarrow\sum_{k\in A}\mu_{b,k}$ $\triangleright$ Set primal variable
7:  $y_{b,A}\leftarrow w_{k^{*}}/\mu_{b,{k^{*}}}$ $\triangleright$ Set dual variable
8:  $w_{k}\leftarrow w_{k}-w_{k^{*}}\mu_{b,k}/\mu_{b,{k^{*}}}$ for all $k\in A$ $\triangleright$ Update weights
9:  $\pi(t)\leftarrow k^{*}$ $\triangleright$ Set priority of coflow $k^{*}$
10:  $A\leftarrow A\setminus\{k^{*}\}$ $\triangleright$ Remove $k^{*}$ from unscheduled coflows
@@
The approximation ratio of Algorithm 1 is established in [Agarwal2018] and formally stated in Theorem 2 below.